A poker hand, consisting of 4 cards, is dealt from a standard deck of 52 cards. Find the probability that the hand contains 4 cards of the same suit.
All help is very much appreciated :)

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

A poker hand, consisting of 4 cards, is dealt from a standard deck of 52 cards. Find the probability that the hand contains 4 cards of the same suit.
-----------
The 1st card is random.
The 2nd is 12/51 = 4/17
Then 11/50
Then 10/49
============
(4/17)*(11/50)*(10/49) = 44/4165
---------
I don't know any poker games where 4 cards is a hand.

Answer by stanbon(75887)

A poker hand, consisting of 4 cards, is dealt from a standard deck of 52 cards. Find the probability that the hand contains 4 cards of the same suit.
------
# of ways to succeed = (13C4)*4
--------
Probability = [13C4 * 4]/[52C4]
------------
Cheers,
Stan H.
------------